Of a random sample of 900 trucks at a bridge, 360 had bad signal lights. Construct a 99 percent confidence interval for the percentage of trucks that had bad signal lights.
Answer #1

Given:

n = 900 trucks, X = 360 had bad signal lights

Find: 99% Confidence interval

= 1 - Confidence level = 1 - 0.99 = 0.01

Calculation:

Critical value:

Z/2 = Z 0.01/2 = 2.58

99% Confidence interval:

(0.36, 0.44)